How would I find the points at which the parabola formed y=x^2 and the line formed by y=2x+3 intersect?
Answer by nerdybill(7384)

How would I find the points at which the parabola formed y=x^2 and the line formed by y=2x+3 intersect?
.
Set them equal to each other and solve for x:
x^2 = 2x + 3
x^2 -2x = 3
x^2 -2x -3 = 0
(x-3)(x+1) = 0
x = {3,-1}
.
To find the corresponding y, plug into either of the two equations and solve for y:
when x=3
y=x^2=3^2 = 9
y=x^2=(-1)^2 = 1
.
Therefore, the points are
(3, 9)
and
(-1, 1)